# New Permutation Decomposition Techniques For Efficient Homomorphic Permutation

### v5

@@ 4.2.3. Complexity Comparison
Table 1 presents existing HMMs under row (or column)-major encoding, where matrices are mapped to the plaintext (vector) space of the batch-encoding HE schemes in a row-wise or column-wise unfolded manner. Jiang et al.’s scheme (Jiang et al., 2018) achieves optimal complexity when considering amortized runtime. Rizomiliotis and Triakosia (Rizomiliotis and Triakosia, 2022) achieve a reduced number of Mults and Rots with a $d$-fold encoding space. However, the increased encoding space may lead to expanded ciphertexts with higher latency per basic operation. Huang et al. (Huang and Zong, 2023) and Zhu et al. (Zhu et al., 2023) design HMMs for rectangular matrices but with modest performance for square matrices. They also require a hypercube packing method supported by BGV-typed HE schemes(Gentry et al., 2012; Brakerski et al., 2014), lacking compatibility with the one-dimensional encoding space. Our method does not require a hypercube structure, so it is compatible to all batch-encoding HE schemes. We also offer configurability in terms of complexity, encoding space, and rotation keys. Notably, we only needs $5\log{n}$ Rots for the maximum redundant encoding space: $(d^{\prime}=d,d_{0}=d)$, and $2d$ Rots for the minimum encoding space: $(d^{\prime}=1,d_{0}=1)$.
@@ 5. New Design for Arbitrary Permutation Decomposition
The Benes network-based decomposition (Gentry et al., 2012) combined with its optimal level-collapsing strategy proposed by Halevi and Shoup (Halevi and Shoup, 2014) may form a decomposition chain with multiplication depth $\log{n}-1$ for any given permutation with length $n$, where each factor contains approximately 6 to 7 non-zero diagonals on average. Although this already offers an approximation to ideal decomposition, we aim to present a new solution outperforming such an approximation in certain scenarios.
Figure 3. Proposed multi-group network design for arbitrary homomorphic permutation (using permutation of length 16 as an example)
@@ 5.1. Overall Design
The network directs entries of the input vector to different paths for processing and gathers them to form the output. For any given permutation, its network is generated as follows.
Initially, the network contains a single standby node at level 00 and group 00, which stores all entries of the input vector. Each entry is marked with an “unsolved” tag, indicating that it has not yet been rotated to its target position. The network is then constructed group by group. For $i_{g}\geq 0$, the $i_{g}$-th group is constructed as follows.
A group starts from the minimum level containing unsolved entries. For level $\ell\geq 0$, we obtain the maximum remaining rotation distance $rot_{max}$ among all unsolved entries at this level. We set $rot$, the rotation step for the rotation node at level $\ell+1$, to be the largest power of two smaller than $rot_{max}$. Then we arrange each unsolved node’s destination in the next level in 3 steps:
@@
Step 2. If $r_{rem}\geq rot$, we assign the entry to a rotation node. So we check whether the $i_{dst}$-th position of the rotation node at level $\ell+1$ in the current group is occupied:
@@
After traversing all unsolved entries, we proceed to the next level by setting $\ell\leftarrow\ell+1$. Otherwise, the $i_{g}$-th group’s construction is completed. We change the tag of entries marked as “unsolved” to “solved” and those marked as “deferred” to “unsolved”, and proceed to the $i_{g}+1$-th group.
Scheme Operation $n$ Number of operations at level $1\rightarrow\log{n}$ $\log{(\#ScalarMult)}$
Ours Rotation $2^{10}$$\{1.0,2.0,3.3,3.7,4.0,4.1,4.3,4.3,4.0,3.8\}$ 31.21
(merged Rescale) $2^{11}$$\{1.0,2.0,3.4,3.9,4.2,4.3,4.3,4.4,4.2,4.1,4.0\}$ 31.39
$2^{12}$$\{1.0,2.0,3.6,4.1,4.2,4.4,4.6,4.8,4.9,4.8,4.7,4.2\}$ 31.53
$2^{13}$$\{1.0,2.0,3.8,4.4,4.7,5.0,5.1,5.2,5.2,5.2,5.1,5.0,4.7\}$ 31.70
$2^{14}$$\{1.0,2.0,4.0,4.7,5.1,5.3,5.4,5.4,5.5,5.6,5.5,5.4,5.3,5.0\}$ 31.83
Benes network Rotation $2^{10}$$\{5.6,6.6,6.6,6.6,6.6,6.6,6.5,6.5,6.5,0\}$ 32.10
(Gentry et al., 2012; Halevi and Shoup, 2014) (double-hoisted) $2^{11}$$\{6.6,3.0,6.6,6.6,6.6,6.6,6.5,6.5,6.5,6.4,0\}$ 32.05
$2^{12}$$\{5.6,6.6,6.6,6.6,6.6,6.6,6.5,6.5,6.5,6.4,6.4,0\}$ 32.25
$2^{13}$$\{6.6,3.0,6.6,6.6,6.6,6.6,6.5,6.5,6.5,6.4,6.4,6.4,0\}$ 32.18
$2^{14}$$\{5.6,6.6,6.6,6.6,6.6,6.6,6.5,6.5,6.5,6.4,6.4,6.4,6.3,0\}$ 32.34
Table 2. Average number of rotations in different network constructions
@@ 5.2. Technical Detail
We present the key features and techniques of our proposed network construction.
@@ 5.2.1. Multi-group Network Construction
The selection of $rot$ at each level aims to limit both the number of rotations in each group and the total number of rotation keys to at most $logn$. This restriction may cause occupation conflicts when multiple nodes attempt to transfer entries to a single rotation node.
@@ 5.2.2. Reduction of Masks and Depth
Many masks within our network are unnecessary in real homomorphic evaluation. They can be replaced with copy operations without compromising the correctness of the network. For any group with a maximum level index of $\ell$, the unnecessary masks are those on edges reaching standby nodes, except for the edges from level $\ell-2$ to level $\ell-1$. This corresponds to the yellow, blue, and purple edges in Figure 3. Reducing them to copy operations allows us to eliminate a significant number of plaintext-ciphertext multiplications and reduce the multiplication depth of the overall network to at most $\log{n}-1$, which aligns with the requirement of the ideal decomposition.
@@ 5.2.3. Timing of Rescale
Rescale is the most computationally intensive operation following rotation in homomorphic permutation, and its execution timing is carefully managed in our proposed network. Whenever a rotation node requires rescaling its input, we reduce its computational complexity by integrating the rescaling into the final submodule of the rotation: ModDown, which receives a ciphertext $ct=(a^{\prime},b^{\prime})\in R_{PQ_{\ell}}^{2}$ from the preceding submodules and outputs $(\lfloor a^{\prime}\cdot P^{-1}\rceil,\lfloor b^{\prime}\cdot P^{-1}\rceil)\in R_{Q_{\ell}}^{2}$ (Cheon et al., 2019; Bossuat et al., 2021). Rescaling a ciphertext $ct=(c_{0},c_{1})\in R_{Q_{\ell}}^{2}$ can be viewed as an instance of ModDown computing $(c_{0},c_{1})\in R_{Q_{\ell-1}}^{2}\leftarrow(\lfloor c_{0}\cdot q_{\ell}^{-1}\rceil,\lfloor c_{1}\cdot q_{\ell}^{-1}\rceil)$. Thus, for a ciphertext requiring both rotation and rescaling, we can substitute the final step of the rotation with a ModDown from $Q_{\ell}P$ to $Q_{\ell-1}$: $(\lfloor a^{\prime}\cdot(q_{\ell}P)^{-1}\rceil,\lfloor b^{\prime}\cdot(q_{\ell}P)^{-1}\rceil)\mod Q_{\ell-1}\in R_{Q_{\ell-1}}^{2}$, completing both rotation and rescaling simultaneously.
Scheme BSGS $d1$ Time (ms) Keys Time (ms) Keys Time (ms) Keys Time (ms) Keys
Original (Jiang et al., 2018) $32$ 1211 38 1108 38 961 38 896 38
Original (Jiang et al., 2018) $16$ 1469 30 1437 30 1202 30 $1$ 30
Ours $16$ 1018 23 722 21 598 21 - -
Ours $8$ 1380 23 844 17 600 15 569 15
Table 3. HMT benchmarks under dimension $d=128$ and different decomposition depth $\ell$
$d=8$ $d=16$ $d=32$ $d=64$
Scheme $d^{\prime}$ $m$ Amortized Keys $m$ Amortized Keys $m$ Amortized Keys $m$ Amortized Keys
Original (Rizomiliotis and Triakosia, 2022) d 64 29 ms 23 8 418 ms 42 1 6142 ms 77 - - -
Ours d 64 17 ms 21 8 184 ms 28 1 1842 ms 35 - - -
Ours d/2 128 13 ms 18 16 127 ms 25 2 1191 ms 32 - - -
Ours d/4 256 7 ms 15 32 86 ms 22 4 870 ms 29 - - -
Ours d/8 512 7 ms 9 64 76 ms 19 8 767 ms 26 1 7594 ms 33
Ours d/16 - - - 128 72 ms 12 16 712 ms 23 2 7009 ms 30
Ours d/32 - - - - - - 32 694 ms 15 4 6762 ms 27
Ours d/64 - - - - - - - - - 8 6829 ms 18
Table 4. HMM benchmarks: full decomposition depth and constant multiplication depth $2$
@@ 5.3. Complexity Analysis
The complexity for computing random permutations of different length $n$ in the network is demonstrated in Table 2, where we present the number of rotations at each level (from $1$ to $\log{n}$). For comparison, the corresponding information for the Benes network-based scheme (Gentry et al., 2012; Halevi and Shoup, 2014) is also provided. In their case, each level corresponds to one double-hoisted homomorphic linear transformation (Bossuat et al., 2021) (see Appendix D for implementation details). The logarithm of the number of scalar multiplications for all rotations in the network is also presented. This is obtained with $logN=15$, modulus chain $Q$ of $18$ moduli, and $P$ of $3$ moduli. At any level $\ell$, the rotations are applied to ciphertexts in $R_{Q_{17-\ell}}^{2}$.
From Table 2, we can observe the following: (i) In our network, the number of rotations first increases and then decreases as the level index grows. The initial increase implies that new groups are created to continue rotating conflicting nodes, while the subsequent decrease arises from the gradual reduction in the number of entries each group needs to process. (ii) For the same $n$, our network exhibits a lower rotation complexity, and this advantage is more pronounced for smaller values of $n$. For $\log{n}=10\sim 14$, the corresponding advantage is approximately $46\%\sim 30\%$.
@@ 6. Implementation
We implement the algorithms proposed in the previous sections using the full-RNS CKKS HE scheme (Cheon et al., 2019; Kim and Song, 2019) provided by the Lattigo library (Mouchet et al., 2020). Experiments are conducted on a system with an i7-13700K processor running at 3.40 GHz and a memory of 64G. We present benchmarks for homomorphic matrix transposition, homomorphic matrix multiplication, along with its application in oblivious neural network inference, and arbitrary homomorphic permutations. Unless otherwise specified, the homomorphic encryption parameters are fixed at $\log{QP}\approx 880$ and $\log{N}=15$ across all benchmarks. This provides a 128-bit security level (Albrecht et al., 2015, 2021; Bossuat et al., 2024), an encoding space of $N/2$ complex numbers per ciphertext, and a budget of 18 modulus levels, indexed from 00 to $17$. Note that homomorphic operations at lower modulus levels inherently incur lower computational costs. Thus when we compare different algorithms, we ensure their execution concludes at the same ciphertext modulus level.

### v7

@@ 4.2.3. Complexity Comparison
Table 1 presents existing HMMs under row (or column)-major encoding, where matrices are mapped to the plaintext (vector) space of the batch-encoding HE schemes in a row-wise or column-wise unfolded manner. Jiang et al.’s scheme (Jiang et al., 2018) achieves optimal complexity when considering amortized runtime. Rizomiliotis and Triakosia (Rizomiliotis and Triakosia, 2022) achieve a reduced number of Mults and Rots with a $d$-fold encoding space. However, the increased encoding space may lead to expanded ciphertexts with higher latency per basic operation. Huang et al. (Huang and Zong, 2023) and Zhu et al. (Zhu et al., 2023) design HMMs for rectangular matrices but with modest performance for square matrices. They also require a hypercube packing method supported by BGV-typed HE schemes (Gentry et al., 2012; Brakerski et al., 2014), lacking compatibility with the one-dimensional encoding space. Our method does not require a hypercube structure, so it is compatible to all batch-encoding HE schemes. We also offer configurability in terms of complexity, encoding space, and rotation keys. Notably, we only need $5\log{n}$ Rots for the maximum redundant encoding space: $(d^{\prime}=d)$. Further incorporating the proposed replication optimization yields a Rot complexity of $\left(\frac{4d}{d_{0}}-2d^{\prime}+\frac{2d\log d_{0}}{d^{\prime}}+3\log d^{\prime}\right)$ for $d^{\prime}>1$, and $\left(\frac{3d}{d_{0}}+\frac{2d\log d_{0}}{d^{\prime}}+3\log d^{\prime}\right)$ under the minimum encoding space ($d^{\prime}=1$), which reduces to $3d$ Rots for $d_{0}=1$.
@@ 5. New Design for Arbitrary Permutation Decomposition
The Benes network-based decomposition (Gentry et al., 2012) combined with its optimal level-collapsing strategy proposed by Halevi and Shoup (Halevi and Shoup, 2014) may form a decomposition chain with multiplication depth $\log{n}-1$ for any given permutation with length $n$, where each factor contains approximately 6 to 7 non-zero diagonals on average. Although this can be considered as a rough approximation to full-depth ideal decompositions for permutations whose non-zero diagonals fill up the whole permutation matrix, we aim to present a new solution outperforming such an approximation in certain scenarios.
@@ 5.1. Overall Design
Performing a homomorphic permutation via this network structure proceeds as follows. Given a ciphertext holding the vector to be permuted, the ciphertext first enters the network at the standby node on level 0. From there, each outgoing edge of that standby node multiplies the ciphertext by its associated mask and forwards the result to the edge’s destination node. If the destination is a standby node, it continues this process: masking and forwarding the ciphertext along its outgoing edges. If the destination is a rotation node, it first performs a rotation on the received ciphertext before passing it along via its outgoing edge. Consequently, all paths lead to the same destination at the bottom of the network, yielding the permuted ciphertext.
In the following, we focus on the network construction. Given a permutation $p$ and any input vector $\mathbf{v}$, $p$ rearranges the entries of $\mathbf{v}$ to produce a new vector $\mathbf{v}^{\prime}\leftarrow p(\mathbf{v})$. During the construction of the network for $p$, we let $\mathbf{v}$ enter the network (which initially consists of a single node) and continue to propagate through the network as it is incrementally built level by level. Throughout this process, we implicitly track the remaining rotation distance for each entry of $\mathbf{v}$ as it is carried from one node to another. This distance is initialized as $r_{org}\leftarrow(j-i)\bmod{n}$, where $i$ is the entry’s original position in $\mathbf{v}$, and $j$ its target position in $p(\mathbf{v})$.
Initially, the network contains a single standby node at level 0 and group 0, which stores all entries of the input vector. Each entry is marked with an “unsolved” tag, indicating that it has not yet been rotated to its target position. The network is then constructed group by group. For $i_{g}\geq 0$, the $i_{g}$-th group is constructed as follows.
A group starts from the minimum level containing unsolved entries. For level $\ell\geq 0$, we obtain the maximum remaining rotation distance $rot_{max}$ among all unsolved entries at this level. We set rotrot, the rotation step for the rotation node at level $\ell+1$, to be the largest power of two no greater than $rot_{max}$. Then we arrange each unsolved node’s destination in the next level in 3 steps:
@@
Step 2. If $r_{rem}\geq rot$, we assign the entry to a rotation node. So we check whether the $i_{dst}$-th position of the rotation node at level $\ell+1$ in the current group is occupied (by this point, if the rotation node has not been created, create it):
@@
After traversing all unsolved entries, we proceed to the next level by setting $\ell\leftarrow\ell+1$ if there exists any $r^{(i)}_{rem}>0$ for $0\leq i<n$, where $r^{(i)}_{rem}$ denotes the remaining rotation distance of the unsolved entry indexed $i$ in the original $\mathbf{v}$. Otherwise, the $i_{g}$-th group’s construction is completed. We change the tag of entries marked as “unsolved” to “solved” and those marked as “deferred” to “unsolved”, and proceed to the $i_{g}+1$-th group. Appendix D.1 further demonstrates the correctness of our network construction in detail.
Table 2. Average number of rotations in different network constructions ($\log{n}$ keys)
Scheme Operation $n$ Number of operations at network level $1\rightarrow\log{n}$ #ScalarMult
Ours Rotation $2^{10}$ $\{1.0,2.0,3.3,3.7,4.0,4.1,4.3,4.3,4.0,3.8\}$ $2.484\times 10^{9}$
(merged Rescale) $2^{11}$ $\{1.0,2.0,3.4,3.9,4.2,4.3,4.3,4.4,4.2,4.1,4.0\}$ $2.814\times 10^{9}$
$2^{12}$ $\{1.0,2.0,3.6,4.1,4.2,4.4,4.6,4.8,4.9,4.8,4.7,4.2\}$ $3.101\times 10^{9}$
$2^{13}$ $\{1.0,2.0,3.8,4.4,4.7,5.0,5.1,5.2,5.2,5.2,5.1,5.0,4.7\}$ $3.489\times 10^{9}$
$2^{14}$ $\{1.0,2.0,4.0,4.7,5.1,5.3,5.4,5.4,5.5,5.6,5.5,5.4,5.3,5.0\}$ $3.818\times 10^{9}$
Benes network Rotation $2^{10}$ $\{4.6,5.6,5.6,5.6,4.6,5.6,5.5,5.5,5.5,0\}$ $3.791\times 10^{9}$
(Gentry et al., 2012; Halevi and Shoup, 2014) (double-hoisted) $2^{11}$ $\{5.6,3.0,5.6,5.6,5.6,4.6,5.5,5.5,5.5,5.4,0\}$ $3.844\times 10^{9}$
$2^{12}$ $\{4.6,5.6,5.6,5.6,4.6,5.6,5.5,5.5,5.5,5.4,5.4,0\}$ $4.178\times 10^{9}$
$2^{13}$ $\{5.6,3.0,5.6,5.6,5.6,5.6,4.5,5.5,5.5,5.4,5.4,5.4,0\}$ $4.207\times 10^{9}$
$2^{14}$ $\{4.6,5.6,5.6,5.6,5.6,5.6,4.5,5.5,5.5,5.4,5.4,5.4,5.3,0\}$ $4.508\times 10^{9}$
@@ 5.2. Technical Detail
We present the key features of our proposed network construction along with techniques for efficient network evaluation.
@@ 5.2.1. Multi-group Network Construction
The selection of rotrot at each level aims to limit both the number of rotations in each group and the total number of rotation keys to at most $\log n$. This restriction may cause occupation conflicts when multiple nodes attempt to transfer entries to a single rotation node.
@@ 5.2.2. Reduction of Masks and Depth
Many masks within our network are unnecessary in real homomorphic evaluation. They can be replaced with copy operations without compromising the correctness of the network. For any group with a maximum level index of $\ell$, the unnecessary masks are those on edges reaching standby nodes, except for the edges from level $\ell-2$ to level $\ell-1$. This corresponds to the yellow, blue, and purple edges in Figure 4. Reducing them to copy operations allows us to eliminate a significant number of plaintext-ciphertext multiplications and reduce the multiplication depth of the overall network to at most $\log{n}-1$.
@@ 5.2.3. Timing of Rescale
Rescale is the most computationally intensive operation following Rot in our proposed network for homomorphic permutation, and its execution timing is carefully managed during the network evaluation. Whenever a rotation node requires rescaling its input, we reduce its computational complexity by integrating the rescaling into the final submodule of the rotation: ModDown, which receives a ciphertext $ct=(a^{\prime},b^{\prime})\in R_{PQ_{\ell}}^{2}$ from the preceding submodules and outputs $(\lfloor a^{\prime}\cdot P^{-1}\rceil,\lfloor b^{\prime}\cdot P^{-1}\rceil)\in R_{Q_{\ell}}^{2}$ (Cheon et al., 2019; Bossuat et al., 2021). Rescaling a ciphertext $ct=(c_{0},c_{1})\in R_{Q_{\ell}}^{2}$ can be viewed as an instance of ModDown computing $(c_{0},c_{1})\in R_{Q_{\ell-1}}^{2}\leftarrow(\lfloor c_{0}\cdot q_{\ell}^{-1}\rceil,\lfloor c_{1}\cdot q_{\ell}^{-1}\rceil)$. Thus, for a ciphertext requiring both rotation and rescaling, we can substitute the final step of the rotation with a ModDown from $Q_{\ell}P$ to $Q_{\ell-1}$: $(\lfloor a^{\prime}\cdot(q_{\ell}P)^{-1}\rceil,\lfloor b^{\prime}\cdot(q_{\ell}P)^{-1}\rceil)\mod Q_{\ell-1}\in R_{Q_{\ell-1}}^{2}$, completing both rotation and rescaling simultaneously. This merging strategy is tailored to our network architecture and has negligible effect on conventional homomorphic permutation evaluation via the HLT algorithm, which performs only one Rescale per HLT.
@@ 5.3. Complexity Analysis
The average complexity for evaluating random permutations of varying lengths $n$ is presented in Table 2, based on 20 samples per length. The table details the number of rotations required at each network level, from $1$ to $\log n$. For comparison, the corresponding information for the Benes network-based scheme (Gentry et al., 2012; Halevi and Shoup, 2014) is also provided. In their case, each level corresponds to one double-hoisted HLT (Bossuat et al., 2021) (see Appendix D.2 for implementation details). The number of scalar multiplications (#ScalarMult) for all rotations in the network is also presented. This is obtained with $\log{N}=15$, modulus chain $Q$ of $18$ moduli, and $P$ of $3$ moduli. At any level $\ell$, the rotations are applied to ciphertexts in $R_{Q_{17-\ell}}^{2}$.
From Table 2, we can observe the following: (i) In our network, the number of rotations first increases and then decreases as the level index grows. The initial increase implies that new groups are created to continue rotating conflicting nodes, while the subsequent decrease arises from the gradual reduction in the number of entries each group needs to process. (ii) For the same $n$, our network exhibits a lower rotation complexity, and this advantage is more pronounced for smaller values of $n$. For $\log{n}=10\sim 14$, the corresponding advantage is approximately $1.52\times\sim 1.18\times$. The variation in the overall number of rotations remains modest in our network, with a slightly increasing standard deviation from $1.6$ to $2.4$ as the permutation length grows.
Scheme $d$ BSGS $d1$ Time (ms) Keys Time (ms) Keys Time (ms) Keys Time (ms) Keys
Original (Jiang et al., 2018) 128 $32$ 1211 38 1108 38 961 38 896 38
Original (Jiang et al., 2018) 128 $16$ 1469 30 1437 30 1202 30 $11$ 30
Ours 128 $32$ 868 35 787 35 - - - -
Ours 128 $16$ 1018 23 722 21 598 21 - -
Ours 128 88 1380 23 844 17 600 15 569 15
Original (Jiang et al., 2018) 256 $64$ 4376 70 4027 70 3507 70 3372 70
Original (Jiang et al., 2018) 256 $32$ 4622 46 4509 46 3735 46 3609 46
Ours 256 $64$ 3352 67 2715 67 - - - -
Ours 256 $32$ 2890 39 2186 37 1710 37 - -
Ours 256 $16$ 3766 31 2311 25 1604 23 1485 23
Table 3. HMT benchmarks under dimension $d\in\{128,256\}$ and different decomposition depth $\ell$
$d=8$ $d=16$ $d=32$ $d=64$
Scheme $d^{\prime}$ $m$ Time (ms) Keys $m$ Time (ms) Keys $m$ Time (ms) Keys $m$ Time (ms) Keys
(Jiang et al., 2018; Ma et al., 2024) 1 512 1802 10 128 3338 16 32 4973 28 8 9043 45
Original (Rizomiliotis and Triakosia, 2022) d 64 1697 23 8 3072 42 1 5530 77 - - -
Ours d 64 1046 21 8 1336 28 1 1705 35 - - -
Ours d/2 128 1257 18 16 1730 25 2 2175 32 - - -
Ours d/4 256 1800 15 32 2584 22 4 3283 29 - - -
Ours d/8 512 3354 9 64 4507 19 8 5736 26 1 7050 33
Ours d/16 - - - 128 8757 12 16 11078 23 2 12943 30
Ours d/32 - - - - - - 32 21773 15 4 25720 27
Ours d/64 - - - - - - - - - 8 51616 18
Table 4. HMM benchmarks under full decomposition depth and constant multiplication depth
@@ 6. Implementation
We implement the algorithms proposed in the previous sections using the full-RNS CKKS HE scheme (Cheon et al., 2019; Kim and Song, 2019) provided by the Lattigo library (Mouchet et al., 2020).Experiments are conducted on a system with an i7-13700K processor running at 3.40 GHz and a memory of 64G. We present benchmarks for homomorphic matrix transposition, homomorphic matrix multiplication, along with its application in oblivious neural network inference, and arbitrary homomorphic permutations. Unless otherwise specified, the homomorphic encryption parameters are fixed at $\log{QP}\approx 880$ and $\log{N}=15$ across all benchmarks. This provides a 128-bit security level (Albrecht et al., 2015, 2021; Bossuat et al., 2024), an encoding space of $N/2$ complex numbers per ciphertext, and a budget of 18 modulus levels, indexed from 0 to $17$. When comparing different algorithms, we ensure that their execution concludes at the same ciphertext modulus level.